OpenStudy (anonymous):

A real estate broker sold your house for $72,000. The broker's commission was 6% of the selling price. How much would you get for the house after the commission is paid?

OpenStudy (anonymous):

do you know how much is 6% of the selling price ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

6/100 x 72000 = 4320 dollars what you get is 94 percent can you get the answer ?

OpenStudy (anonymous):

no i get it this the choice A. $76,320 B. $73,200 C. $70,800 D. $67,680

OpenStudy (anonymous):

94 percent = 100 percent - 6 percent you know the 100 percent, which is 72000 dollars and you also know the 6 percent, which is 4320 dollars 94 percent = 72000 - 4320 = 67680 dollars answer is D

OpenStudy (anonymous):

ooo i get it now thank you

OpenStudy (anonymous):

sure ^^